# Symmetrizing the Constraints – Density Matrix Renormalization Group for Constrained Lattice Models

### v1

@@ III.2 Ground state phase diagram
We show the obtained dimer-dimer correlation function $S(\mathbf{k})=\frac{1}{N}\sum_{i,j}e^{-i\mathbf{k}\cdot(\mathbf{r}_{i}-\mathbf{r}_{j})}\langle\psi\rvert(n_{i}-\frac{1}{6})(n_{j}-\frac{1}{6})\lvert\psi\rangle$ in the Brillouin zone at $V=-1$ (inside the columnar phase), $V=-0.2$ (inside the VBS phase) and $V=0.9$ (inside the QSL phase). The OBC along $x$-direction breaks the translational symmetry, rendering $k_{x}$ not a good quantum number. However, our computation of $S(\mathbf{k})$ focused solely on the correlations in the bulk region, which is sufficiently far from the boundary, where the translational symmetry remains largely intact. In Fig. 4 (a) and (b), it is clear that the Bragg peaks are at one of three $M$ points, as we chose only bonds towards the top left direction in this calculation. As $V$ moves to $-0.2$ (Fig. 4 (a) and (c)), the Bragg peak moves to the $X$ point, which according to the previous literature [22], is the wave vector of the $\sqrt{12}\times\sqrt{12}$ order from dimer correlation function. Here, again, we observe one of the three pairs of the $X$ peaks, and the other two can be seen if one looks at bonds in other directions. It is interesting to observe that inside the VBS phase, the Bragg peak at the $M$ point is also present, although it is secondary. This is also consistent with the results in previous QMC works [22, 23, 24]. As $V$ moves inside the QSL phase (Fig. 4 (a) and (d)), the dimer correlation in the entire BZ flattens, as there is no long-range order in the dimer arrangement and both dimer and vison spectra are gapped [24]. The system acquires with $\mathbb{Z}_{2}$ topological order and one can detect such long-range entanglement with the scaling of the entanglement entropy [40, 41, 45, 51], as we now turn to.
@@ III.3 Entanglement entropy inside the $\mathbb{Z}_{2}$ QSL phase
Moreover, on a finite size system, the EE can also reveal the degeneracy of the ordered ground state. For example, it is expected on a finite size lattice
$S=al+d,$ (5)
where $d$ is the logarithm of the number of ground state degeneracy due to the symmetry breaking [72]. One can also test this behavior on the cylinder geometry inside the columnar phase, where $d=\ln(2)$ is expected as only one of the three columnar directions is favored.
In our DMRG simulation of cylindrical geometry, the spin liquid phase is two-fold degenerated, one from each topological sector ($M=\pm 1$). Previous studies revealed that the constant term in the scaling of EE differs within the two-dimensional ground state eigen-space when the bipartition boundary is non-contractable [43, 42, 45]. Luckily for DMRG, the computation finds the optimized state with minimal EE (maximal $\gamma$). And such minimal entropy state (MES) is a combination of states from two sectors, giving $\langle M\rangle=0$. Namely, one needs to perform DMRG without separating the two topological sectors.
Figure 5: Extrapolation of EE in $\mathbb{Z}_{2}$ quantum spin liquid and columnar phases. With a cylinder geometry and smooth boundary, the TEE (the intercept of the dashed fitting line) extrapolates to $-\ln(2)$ using data with boundary length $l=L_{y}=2,4,6$, not only at the RK point with $V=1$ but also inside the QSL phase with $V=0.9,0.92,0.94,0.96$ and 0.98. Inside the columar phase with $V=-1$, EE fluctuates around $\ln(2)$, coming from the ground state degeneracy of the symmetry-breaking pattern of columnar order on a cylinder. The inset shows an example of a smooth cut, where the red bonds are considered as subregion $A$ and black bonds as $B$.
Fig. 5 shows the result of von-Neumann entanglement entropy $S^{\textrm{vN}}$ for a few $V$, where the system is believed to be in the $\mathbb{Z}_{2}$ spin liquid phase. We reverted to the smooth boundary in this computation since we do not need to satisfy the VBS phase here. We compute the EE with $L_{y}=2,4,6$ and $L_{x}$ is kept to be 24, where EE has stabilized and can reflect bulk information. The system is cut in half along the y direction, creating a smooth boundary with length $l=L_{y}$. Using only the three smallest possible circumferences, we are able to obtain intersections around 0.7(1) by linear fitting, which is close to the anticipated $\gamma=\ln(2)\approx 0.693$. We note that such precise determination of the TEE at the thermodynamic limit, inside the $\mathbb{Z}_{2}$ QSL phase away from the RK point, is reported for the first time being reported.
We also perform the DMRG in the columnar phase ($V=-1$), and the resulting EE is around $\ln(2)$ as the boundary length increases. This indicates a nearly separable state with only $d=\ln(2)$ contribution from the ground state degeneracy.
@@ IV Discussion
The idea of symmetrizing constraint and the associated dynamic storage of quantum numbers can be readily applied on other constrained models, such as PXP model [76, 77], which also describes synthetic quantum simulators. Our method can also serve as a useful tool to detect the topological double semion order in kagome dimer model with an extra term [78], which QMC would face the sign problem.
@@ Acknowledgements.
We thank Menghan Song, Hongyu Lu, and Min Long for valuable discussions on the DMRG implementation. We thank Fabien Alet, Sylvain Capponi and Junchen Rong for the inspiring discussion on the physics of constrained lattice models and we are in particular grateful for the insightful suggestions from Fabien Alet on the narrative and presentation of the manuscript. TTW, XXR and ZYM acknowledge the support from the Research Grants Council (RGC) of Hong Kong (Project Nos. 17301721, AoE/P-701/20, 17309822, HKU C7037-22GF, 17302223, 17301924), the ANR/RGC Joint Research Scheme sponsored by RGC of Hong Kong and French National Research Agency (Project No. A_HKU703/22). We thank the HPC2021 system under the Information Technology Services, The University of Hong Kong, as well as the Beijing PARATERA Tech CO., Ltd. (URL: https://cloud.paratera.com) for providing HPC resources that have contributed to the research results reported in this paper.

### v3

@@ III.2 Ground state phase diagram
We show the obtained dimer-dimer correlation function $S(\bm{k})=\frac{1}{N^{2}}\sum_{i,j}e^{-i\bm{k}\cdot(\bm{r}_{i}-\bm{r}_{j})}\langle\psi\rvert(\hat{n}_{i}-\frac{1}{6})(\hat{n}_{j}-\frac{1}{6})\lvert\psi\rangle$ in the Brillouin zone at $V=-1$ (inside the columnar phase), $V=-0.2$ (inside the VBS phase) and $V=0.9$ (inside the QSL phase). The OBC along $x$-direction breaks the translational symmetry, rendering $k_{x}$ not a good quantum number. However, our computation of $S(\bm{k})$ focused solely on the correlations in the bulk region, which is sufficiently far from the boundary, where the translational symmetry remains largely intact. In Fig. 4 (a) and (b), it is clear that the Bragg peaks are at one of three $\bm{M}$ points, as we chose only bonds towards the top left direction in this calculation. As $V$ moves to $-0.2$ (Fig. 4 (a) and (c)), the Bragg peak moves to the $\bm{X}$ point, which according to the previous literature [22], is the wave vector of the $\sqrt{12}\times\sqrt{12}$ order from dimer correlation function. Here, again, we observe one of the three pairs of the $\bm{X}$ peaks, and the other two can be seen if one looks at bonds in other directions. It is interesting to observe that inside the VBS phase, the Bragg peak at the $\bm{M}$ point is also present, although it is secondary. This is also consistent with the results in previous QMC works [22, 23, 24]. As $V$ moves inside the QSL phase (Fig. 4 (a) and (d)), the dimer correlation in the entire BZ flattens, as there is no long-range order in the dimer arrangement and both dimer and vison spectra are gapped [24]. The system acquires with $\mathbb{Z}_{2}$ topological order and one can detect such long-range entanglement with the scaling of the entanglement entropy [40, 41, 45, 51], as we now turn to. The results of the order parameters and entanglement entropy scanning through the ground state phase diagram can be found in Appendix B.
@@ III.3 Entanglement entropy inside the $\mathbb{Z}_{2}$ QSL phase
In our DMRG simulation of cylindrical geometry, the spin liquid phase is two-fold degenerate, one from each topological sector ($M=\pm 1$). Previous studies revealed that the constant term in the scaling of EE differs within the two-dimensional ground state eigen-space when the bipartition boundary is non-contractable [43, 42, 45]. Luckily for DMRG, the computation finds the optimized state with minimal EE (maximal $\gamma$). And such minimal entropy state (MES) is a combination of states from two sectors, giving $\langle M\rangle=0$. Namely, one needs to perform DMRG without separating the two topological sectors.
Figure 5: Extrapolation of EE in $\mathbb{Z}_{2}$ quantum spin liquid and columnar phases. With a cylinder geometry and smooth boundary, the TEE (the intercept of the dashed fitting line) extrapolates to $-\ln(2)$ using data with boundary length $l=L_{y}=2,4,6$, not only at the RK point with $V=1$ but also inside the QSL phase with $V=0.9,0.92,0.94,0.96$ and 0.98. The inset shows an example of a smooth cut, where the red bonds are considered as subregion $A$ and black bonds as $B$.
Fig. 5 shows the result of von-Neumann entanglement entropy $S^{\textrm{vN}}$ for a few $V$, where the system is believed to be in the $\mathbb{Z}_{2}$ spin liquid phase. We reverted to the smooth boundary in this computation since we do not need to satisfy the VBS phase here. We compute the EE with $L_{y}=2,4,6$ and $L_{x}$ is kept to be 24, where EE has stabilized and can reflect bulk information. The system is cut in half along the y direction, creating a smooth boundary with length $l=L_{y}$. Using only the three smallest possible circumferences, we are able to obtain intersections around 0.7(1) by linear fitting, which is close to the anticipated $\gamma=\ln(2)\approx 0.693$. We note that such precise determination of the TEE at the thermodynamic limit, inside the $\mathbb{Z}_{2}$ QSL phase away from the RK point, is reported for the first time.
@@ IV Discussion
The idea of symmetrizing constraint and the associated dynamic storage of quantum numbers can be readily applied on other constrained models, such as PXP model [80, 81], which also describes synthetic quantum simulators. Our method can also serve as a useful tool to detect the topological double semion order in kagome dimer model with an extra term [82], where QMC would face the sign problem.
@@ Acknowledgements.
We thank Menghan Song, Hongyu Lu, and Min Long for valuable discussions on the DMRG implementation. We thank Fabien Alet, Sylvain Capponi, and Junchen Rong for the inspiring discussion on the physics of constrained lattice models, and we are in particular grateful for the insightful suggestions from Fabien Alet on the narrative and presentation of the manuscript. TTW, XXR and ZYM acknowledge the support from the Research Grants Council (RGC) of Hong Kong (Project Nos. 17301721, AoE/P-701/20, 17309822, HKU C7037-22GF, 17302223, 17301924), the ANR/RGC Joint Research Scheme sponsored by RGC of Hong Kong and French National Research Agency (Project No. A_HKU703/22). We thank the HPC2021 system under the Information Technology Services, The University of Hong Kong, as well as the Beijing PARATERA Tech CO., Ltd. (URL: https://cloud.paratera.com) for providing HPC resources that have contributed to the research results reported in this paper.
Appendix A Implementation of local symmetries in MPO and MPS
The total Hilbert space is defined as $\mathscr{H}=\bigotimes_{b}\mathscr{H}_{b}$, where each local Hilbert space on bond $b$ is a direct sum of empty and occupied dimer states, i.e., $\mathscr{H}_{b}=\lvert 0\rangle_{b}\oplus\lvert 1\rangle_{b}$. The conservation of the numbers of dimers can be viewed of as $N_{\textrm{site}}$ independent $\mathrm{U}(1)$ symmetries. Let $\bm{\theta}=(\theta_{1},\cdots,\theta_{N_{\textrm{site}}})$ and $\hat{\bm{N}}=(\hat{n}_{1},\cdots,\hat{n}_{N_{\textrm{site}}})$ be two $N_{\textrm{site}}$-dimensional vectors, where each $\theta_{v}$ is an angle from 0 to $2\pi$, and operator $\hat{n}_{v}=\sum_{b\in v}\hat{n}_{b}$ counts the number of dimers on bond connected to site $v$. The basis states of local Hilbert space $\mathscr{H}_{b}$ on bond $b$ transform under the $\mathrm{U}(1)$ rotations as,
$e^{i\bm{\theta}\cdot\hat{\bm{N}}}\lvert 0\rangle_{b}=\lvert 0\rangle_{b}\text{, and }$
$e^{i\bm{\theta}\cdot\hat{\bm{N}}}\lvert 1\rangle_{b}=e^{i\bm{\theta}\cdot\bm{\mu}_{b}}\lvert 1\rangle_{b},$
where elements in vectors $\bm{\mu}_{b}=(0,\cdots,1,0,\cdots,1,0,\cdots,0)$ equal to 1 for the two sites $v$ connected to bond $b$ and 0 otherwise, i.e.,
$[\bm{\mu}_{b}]_{v}=\begin{cases}1,&\text{if }b\in v\\ 0,&\text{otherwise}\end{cases}$
For QDM, the constrained Hilbert space $\mathscr{H}_{\textrm{QDM}}\subset\mathscr{H}$ contains wave functions that are superpositions of product states under the one dimer per site constraint,
$\lvert\Psi\rangle=\sum_{\{\sigma_{b}=0,1\}}a_{\{\sigma_{b}\}}\bigotimes_{b}\lvert\sigma_{b}\rangle,$
summing over states where $\sum_{b}\sigma_{b}\bm{\mu}_{b}=(1,1,\cdots,1)$.
The Hamiltonian commutes with the transformation $e^{i\bm{\theta}\cdot\hat{\bm{N}}}$. One has
$\langle\{\sigma_{b}\}\rvert H\lvert\{\sigma^{\prime}_{b}\}\rangle\neq 0\text{,only if }\sum_{b}(\sigma_{b}-\sigma^{\prime}_{b})\bm{\mu}_{b}=(0,0,\cdots,0)$
Therefore, we can track these quantum numbers as $N_{\textrm{site}}$-dimensional vector for MPO and MPS.
Figure 6: Schematics for local tensors Local tensors in (a) MPS and (b) MPO. The arrows indicate the flows of quantum numbers (vector) $\bm{q}$. The vertical (horizontal) arrows represent the physical (virtual) spaces. The quantum numbers add up respectively when the spaces are fused.
Fig. 6 shows schematics for local tensors. Following the directions of the arrows in (a) and (b), one can mark each dimension with vector so that nonzero elements in each local tensor of an MPS (MPO) locate only in blocks where $\bm{q}_{1}+\bm{q}_{3}=\bm{q}_{2}\,(+\bm{q}_{4})$, where the quantum numbers at the physical dimensions is defined as $\sigma_{b}\bm{\mu}_{b}$.
Appendix B Ground state result
Figure 7: Observables measured from MPS. Result of dimmer correlation at (a) $\bm{X}$ and (b) $\bm{M}$ points, and (c) von-Neumann entanglement entropy. The three panels share the same legend on the top.
In this section, we present observables obtained from DMRG across the phase diagram.
The dimmer-dimmer correlations $D(\bm{X})$ and $D(\bm{M})$ are measured using the resulting MPS with the definition mentioned in the main text, except the $D(\bm{M})$ are computed for horizontal dimers, which is favored in the DMRG computation for negative $V$.
As shown in Fig. 7 (a) & (b), $D(\bm{X})$ ($D(\bm{M})$) is the peak in the Brillouin zone, serving as an order parameter for the VBS (columinar) phase. At around $V=-1$, there is a jump for both order parameters, which suggests a first-order transition between the two phases.
The entanglement entropy (EE) exhibits a sharp increase following the first-order transition and continues to rise with increasing $V$.
Figure 8: Entanglement entropy and topological sector expectation value. Result of (a) entanglement entropy and (b) the expectation value of operator $M$ defining the two topological sectors for $V$ from 0.5 to 1. The legend is shown on the top of Fig. 7.
Fig. 8 presents a zoomed-in view of the EE and the expectation value of $M$ in the vicinity of the spin liquid phase. For small values of $V$, the DMRG algorithm correctly identifies the ground state within the appropriate topological sector of the VBS phase: specifically, $M=+1$ ($-1$) when $L_{x}/2$ is even (odd). Within the spin liquid phase, the two topological sectors become nearly degenerate. For states deep inside this phase, the DMRG algorithm tends to converge to a superposition of the two sectors, resulting in an average value of $\langle M\rangle$ that deviates from ±$\pm 1$ to around 0, that is the MES mentioned in the main text [43, 42, 45], accompanied by a characteristic reduction in EE.